### v2

# Pulsed radio emission from a Central Compact Object

Abstract
The high magnetic fields and rapid spins of young pulsars associated with supernova remnants, such as the Crab and the Vela, established the standard pulsar model in which massive stellar explosions produce rapidly rotating, radio-luminous neutron stars. Central Compact Objects (CCOs), identified in X-rays at the centers of other remnants, challenged this view, as decades of searches yielded no radio detections. Here we show that the prototypical young CCO 1E 1207.4–5209 is in fact a faint radio pulsar rotating at the 0.4s X-ray period. Analysis of its polarization indicates that the radio beam intersects our line of sight near the magnetic pole, affirming its radio faintness’ being intrinsic. Once its supernova remnant dissipates, this source would be misidentified as an apparently gigayear-old pulsar. The CCO’s low radio flux density may explain why many supernova remnants lack detectable radio pulsars and suggests a hidden population of young, slowly rotating neutron stars.
{affiliations}
State Key Laboratory of Radio Astronomy and Technology, National Astronomical Observatories, Chinese Academy of Sciences, Beijing 100101, China
Centre for Astrophysics and Supercomputing, Swinburne University of Technology, Hawthorn, VIC 3122, Australia
Fakultät für Physik, Universität Bielefeld, Postfach 100131, Bielefeld D-33501, Germany
New Cornerstone Science Laboratory, Department of Astronomy, Tsinghua University, Beijing 100084, China
School of Arts and Sciences, Qingdao Binhai University, Huangdao District, Qingdao 266555, China
South African Radio Astronomy Observatory, Liesbeek House, River Park, Gloucester Road, Cape Town, 7700, South Africa
Department of Physics and Astronomy, Haverford College, 370 Lancaster Avenue, Haverford, PA 19041, USA
School of Astronomy and Space Science, Nanjing University, Nanjing 210023, China
Key Laboratory of Modern Astronomy and Astrophysics, Nanjing University, Ministry of Education, Nanjing 210023, China
Department of Physics, CCIS 4-183, University of Alberta, Edmonton, AB T6G 2E1, Canada
Young neutron stars manifest themselves in diverse forms, including rotation-powered pulsars, magnetars, and the enigmatic class of central compact objects (CCOs). CCOs have previously been defined as point-like X-ray sources at the geometrical centres of supernova remnants (SNRs), without optical counterparts, radio emission, or associated pulsar wind nebulae (PWNe) [1].
Ten members of this class are currently confirmed***http://www.iasf-milano.inaf.it/~deluca/cco/main.htm. Among them, 1E 1207.4-5209 in the SNR PKS 1209-51/52 (G296.5+10.0) is archetype. The associated SNR has an estimated age of $\sim$2–20 kyr [2, 3] and a distance of just $\sim$1.4 kpc [3]. It exhibits 424-ms X-ray pulses [4], confirming its neutron-star nature, and unusual absorption features in its spectrum [5, 6], interpreted as cyclotron or atomic lines. The agreement of the dipolar B field ($\sim 8\times 10^{10}$ G) inferred from the cyclotron line interpretation [6], and the dipolar B field ($\sim 9.8\times 10^{10}$ G) inferred from its period and period derivative [7, 8], established 1E 1207.4-5209 as a low magnetic field neutron star.
Over the past two decades, searches for radio pulses from CCOs have yielded only upper limits [9, 10, 11]. This persistent radio silence—alongside their purely thermal X-ray emission, exceptionally low spin-down rates, and absence of counterparts at other wavelengths—has led to the hypothesis that CCOs may represent a distinct evolutionary pathway among young neutron stars [12]. One proposed explanation is the “anti-magnetar” model [13, 8], in which CCOs are born with intrinsically weak external dipole fields, insufficient to sustain active magnetospheres. An alternative scenario involves magnetic-field burial by fallback accretion [14, 15], where initially strong fields are submerged beneath the crust by a large amount of mass accreted soon after neutron star formation [16, 17, 18, 19] and gradually re-emerge over timescales of $10^{3}$$10^{5}$ years. A third possibility invokes a stochastic dynamo during the proto-neutron star phase [20, 21], which generates a highly tangled internal field, with either a weak or absent dipole component.
[…]
To further characterize its emission, we obtained an 8.5-hour dual-frequency MeerKAT observation in split-array mode at UHF and L-band. The pulse flux varied significantly over time and frequency (Methods and Extended Data Fig. 2), consistent with interstellar scintillation typical of pulsars at that DM  [24], which may have contributed to previous non-detections. Figure 1 shows the integrated polarization profiles from both bands, and Table 1 summarizes key radio and X-ray parameters.
The radio properties of PSR J1210-5226 provide a unique opportunity to assess how CCOs relate to the broader pulsar population. The measured profile width $W_{10}$ of $\sim$17% is relatively wide but within the observed range for ordinary pulsars with similar spin periods [25]. Its linear polarization fraction however ($\sim$77%) is unusually high for pulsars with such low spin-down energy, and is more typical of young, high-$\dot{E}$ systems [26], suggesting a dynamically young or recently reconfigured magnetosphere. Since the SNR that is associated with and surrounds the pulsar is $\sim$14 kyr [3] old, the pulsar’s characteristic age of $\sim$300 Myr derived from its X-ray timing parameters suggests that the initial spin period was very close to what we observe today.
Additional constraints on the pulsar geometry are provided by a Rotating Vector Model (RVM) fit to the polarization position angle swing. Although the data yield relatively weak constraints due to the limited signal-to-noise ratio and position angle phase coverage, the best-fit solution favors a magnetic inclination angle ($\alpha$) between $\sim$17 and 31, and an impact parameter ($\beta$) between $\sim-7^{\circ}$ and $11^{\circ}$, implying a small viewing angle ($\zeta)\sim 15^{\circ}$ between the rotation axis and our line of sight (Methods and Extended Data Figure 3). Such a geometry is consistent with a relatively aligned rotator and explains the simple, nearly sinusoidal X-ray pulse profile of PSR J1210-5226, which is thought to arise from thermal emission from a single hot spot always in view [4, 6, 8]. The lack of narrow features in the X-ray profile may support a low magnetic field interpretation, although the field configuration remains uncertain. Taken together, the emission profile, polarization properties, and inferred geometry suggest that PSR J1210-5226, while radio faint, exhibits characteristics broadly consistent with ordinary pulsars [27] and may represent a previously unrecognized evolutionary link between CCOs and the radio-emitting pulsar population.
1E 1207.45209-5209 is located within the bilaterally symmetric supernova remnant PKS 1209-51/52 [22]. Radio polarization observations reveal a distinct east-west RM asymmetry in the remnant, with mean values of +28 rad m-2on the eastern limb and 14-14 rad m-2 on the western limb. The foreground RM along the line of sight to the SNR is estimated to lie in the range of +10 to +20 rad m-2 [28]. The RM value associated with the CCO (++10 rad m-2) is consistent with the range toward the supernova remnant, albeit with considerable uncertainty for the latter measurements. This consistency provides no evidence of a strong magneto-ionic environment surrounding the CCO, such as one that could be produced by a magnetized fallback disk. This agrees with HST and VLT measurements, which suggest an ultra-low mass ($10^{-10}~M_{\odot}$) for any potential fallback disk surrounding this CCO [29].
The discovery of radio pulsations from 1E 1207.4-5209 confirms that at least some CCOs possess active magnetospheres capable of sustaining coherent radio emission. Previous searches for radio pulsations from this source and other CCOs using sensitive instruments such as the GBT, FAST, and earlier MeerKAT campaigns had returned null results [10, 11, 23], reinforcing the perception that CCOs are intrinsically radio-quiet. Our detection of PSR J1210-5226 establishes the first radio CCO, offering a direct observational link between this class and the broader pulsar population. While our MeerKAT campaign yielded non-detections for the other three CCO targets (Methods), these results are not necessarily evidence for radio silence across the population. Figure 2 shows the corresponding flux density upper limits. They may reflect limited beaming fractions [30, 31], interstellar scintillation, low intrinsic luminosity, or limited instrumental sensitivity [32, 33]. It is also possible that 1E 1207.4-5209 may represent a rare transitional phase in which re-emerging magnetic fields briefly enable coherent radio emission. Another possibility is that glitches, observed in 1E 1207.4-5209 in 2015[34, 35], caused by crustquakes may change the magnetic inclination angle or affect the geometry of the magnetic field lines in the vicinity of the polar cap, leading to temporary radio emission.
Our detection further confirms that 1E 1207.4-5209 hosts a magnetic field, relatively weak compared to the broader pulsar population, yet strong enough to produce detectable coherent radio emission. This, together with its high bolometric X-ray luminosity [4], challenges the interpretation of CCOs as objects that are intrinsically radio inactive. Two broad explanations may account for this detection. One involves magnetic-field topology: a complex internal field configuration with highly multipolar structure, rather than a uniformly weak dipole field, could reconcile the properties of 1E 1207.4-5209 with the persistent radio silence of other CCOs. As discussed in the Methods, a stochastic dynamo during the proto-neutron star phase may generate a tangled magnetic field residing within the crust, the evolution of which allows for localized multipolar magnetic structures, providing conditions favorable for radio emission. The second possibility invokes magnetic-field evolution provoked by fallback accretion: 1E 1207.4-5209, with its longer spin period and possibly older age compared to the other two X-ray pulsating CCOs, may have allowed sufficient time for a previously buried dipole field [16, 19] to re-emerge [17, 18, 14, 36], potentially accelerated by the 2015 glitch [34, 35, 37]. Distinguishing between these scenarios remains difficult, as the timescales for magnetic restructuring are long compared to existing observations. Continued radio monitoring, particularly of spin-down rate variations, along with deeper searches for additional radio-loud CCOs, will be essential for clarifying the physical mechanisms and evolutionary pathways linking CCOs to the broader neutron star population.
Once the supernova remnant has dissipated, PSR J1210–5226 may well continue to be an active radio pulsar for a Gyr or more and be indistinguishable from the normal pulsar population, albeit of low luminosity. It gives support to early suggestions[38] that some or indeed a high fraction of neutron star pulsars are “injected” into the population with relatively long spin periods near 0.5 s. PSR J1210–5226’s low luminosity implies that a potentially large population of similar objects may inhabit the galaxy and explains the absence of detectable radio emission from other putative neutron stars in binaries and other SNRs.
This discovery firmly establishes 1E 1207.4-5209 as a radio-emitting neutron star, demonstrating that CCOs are not an entirely radio-silent population. It provides the first direct observational link between these enigmatic thermal X-ray sources and ordinary pulsars. Continued monitoring of PSR J1210-5226, together with deeper searches for radio emission from other CCOs, will be crucial to determine whether radio activity is a general but elusive property of the class or a unique trait of this object. These results underscore the importance of high-sensitivity surveys with instruments like MeerKAT and highlight the potential of next-generation facilities such as the Square Kilometre Array to uncover additional radio-active CCOs within the Galactic neutron star population.
Table 1: Parameters for PSR J1210-5226
Parameter Value
Parameters from MeerKAT (544–1088 MHz)
Right Ascension (J2000)a $12^{\rm h}\;10^{\rm m}\;00^{\rm s}.91$
Declination (J2000)a $52^{\circ}\;26^{{}^{\prime}}\;28^{{}^{\prime\prime}}.4$
Period, $P$ (s) 0.4241307(1)
Predicted $P$ from X-ray timingb (s) 0.4241307605(1)
Dispersion measure, DM (pc cm-3) 69.0(4)
Rotation Measure, RM (rad m-2) ++10(1)
Epoch of ephemeris (MJD) 60314.961
Pulse width at 50% of peak, $W_{50}$ () 33
Pulse width at 10% of peak, $W_{10}$ () 61
[…]
Percentage linear polarization ($L/I$) 77(5)
Percentage circular polarization ($V/I$) ++30(4)
Percentage absolute circular polarization ($\left|V\right|/I$) 32(4)
Parameters from X-ray observations[37]
Right Ascension (J2000) $12^{\rm h}\;10^{\rm m}\;00^{\rm s}.9126(29)$
Declination (J2000) $52^{\circ}\;26^{{}^{\prime}}\;28^{{}^{\prime\prime}}.303(42)$
Period, $P$ (s)c 0.424130756710(26)
Period derivative, $\dot{P}$ (s s-1)c $2.015(49)\times 10^{-17}$
Epoch of ephemeris (MJD) 58144.000
Surface dipole magnetic field, $B_{s}$ (G)d $9.8\times 10^{10}$
Spin-down luminosity, $\dot{E}$ (erg s-1)d $1.1\times 10^{31}$
Characteristic age, $\tau_{c}$ (Myr)d 303
Distance Estimates
From SNR PKS 1209-51/52 (kpc)[3] 1.4
From DM: NE2001 / YMW16 (kpc)[39, 40] 1.9 / 1.5
Note. The radio period is consistent, within uncertainties, with the period extrapolated from the X-ray timing
ephemeris[37], confirming the identification of the radio and X-ray sources. The Stokes parameters were defined
using the PSR/IEEE convention described in [41].
a The telescope was pointed to the X-ray coordinates, whose position is known to better than $1^{{}^{\prime\prime}}$, well within
the MeerKAT UHF tied-array beam (FWHM $\approx 0.5^{{}^{\prime}}$).
b Predicted using X-ray ephemeris parameters ($P$, $\dot{P}$) at MJD 58144 [37].
c From the post-glitch timing solution (2016-2023).
d Derived parameters ($B_{s}$, $\dot{E}$, $\tau_{c}$) are based on the pre-glitch timing solution (2002-2014). The characteristic
age is not indicative of the true age, which is estimated to be $\sim$14 kyr of its host SNR, PKS 1209-51/52[3].
Figure 1: Polarization profiles of PSR J1210-5226. Integrated polarization profiles from two MeerKAT observations: a 4-hour UHF-band observation (544–1088 MHz) on 2024 January 5 (bottom), and the first 4.25-hour segment of the L-band observation (856–1712 MHz) on 2025 October 18 (top). In each panel, black, red, and blue lines represent the total intensity, linear polarization, and circular polarization, respectively. Black points in the upper panels indicate the polarization position angle (PA) of the linear component.
Figure 2: Sensitivity curves for the four CCOs observed with a 4-hour MeerKAT integration in the UHF band (544–1088 MHz). Left: Theoretical upper limits on radio flux density for the two CCOs with known X-ray spin periods, plotted as a function of dispersion measure (DM) for various assumed intrinsic duty cycles (d.c.). The red star marks PSR J1210-5226 at its observed DM and measured mean flux density from this work. The predicted DMs from Galactic electron density models are 127 and 82 pc cm-3 (YMW16, NE2001, respectively) for 1E 1207.4-5209, and 157 and 417 pc cm-3 for RX J0822.0-4300. Right: Theoretical flux density limits for the two CCOs without known spin periods, shown as a function of spin period, assuming a 10% intrinsic duty cycle. In all cases, dispersive smearing is the dominant source of pulse broadening and sensitivity degradation.